# On the uncertainty of the White Dwarf Astrophysical Gravitational Wave Background

### v1

@@ Abstract
Context. The astrophysical gravitational wave background (AGWB) is a stochastic GW signal that is emitted by different populations of inspiralling binary systems containing compact objects. In the frequency range between 10-4 and 10-1 Hz it will be detected by future space-based gravitational wave detectors like the Laser Interferometer Space Antenna (LISA). Recently, Staelens & Nelemans (2024) concluded that the white dwarf (WD) contribution to the AGWB dominates over that of black holes (BHs).
@@
Methods. We use the code developed by Staelens & Nelemans (2024) to determine the WD component of the AGWB. We use a metallicity dependent SFRD based on Chruślińska & Nelemans (2019); Chruślińska et al. (2020, 2021) to construct five different SFRD models. We use four different population models that use different common-envelope treatment and six different metallicities for each model.
@@ 1 Introduction
In the visible Universe, there are a very large number of gravitational wave (GW) sources that consist of two stellar mass compact objects, white dwarfs (WDs), neutron stars (NSs) or black holes (BHs). With a GW detector typically only a (tiny) selection of these sources can be individually detected, and many cannot. The Laser Interferometer Space Antenna (LISA) mission (Amaro-Seoane et al., 2017), for instance can individually detect a large number of binaries (mostly double WD) in the Milky Way (e.g. Lipunov et al., 1987; Evans et al., 1987; Hils et al., 1990; Nelemans et al., 2001a; Ruiter et al., 2010; Yu & Jeffery, 2013; Lamberts et al., 2019; Korol et al., 2022; Temmink et al., 2023), a collective stochastic foreground of Galactic sources that is highly anisotropic (e.g. Nissanke et al., 2012; Karnesis et al., 2021) plus a stochastic isotropic signal from all sources in the Universe (e.g. Farmer & Phinney, 2003). This Astrophysical GW Background (AGWB) encodes the combined information about the different source populations (see, e.g. Schneider et al., 2001; Farmer & Phinney, 2003; Kowalska-Leszczynska et al., 2015; Amaro-Seoane et al., 2023). It is a broadband signal (Phinney, 2001) that extends from the high-frequency band covered by the LIGO/Virgo/KAGRA detectors (LIGO Scientific Collaboration et al., 2015; Acernese et al., 2015; Aso et al., 2013) to the mHz band covered by future space GW detectors such as LISA (Amaro-Seoane et al., 2017), Taiji (Luo et al., 2021) or TianQin (Luo et al., 2016).
@@
The AGWB is not only interesting in itself as probe of the large population of binaries across cosmic time that cannot be detected individually, but also competes with potential other stochastic GW backgrounds that have an origin in the Early Universe and probe fundamental physics (see e.g. Caprini et al., 2009; Binétruy et al., 2012; Auclair et al., 2023).
@@
In the LISA band, the AGWB due to BH mergers could be detected (e.g. Dvorkin et al., 2016; Bavera et al., 2022; Babak et al., 2023; Lehoucq et al., 2023), but Staelens & Nelemans (2024) recently showed that the AGWB due to WD binaries (studied earlier by Farmer & Phinney, 2003) likely dominates. However, that study only considered a single model for the double WD population in the Universe and a simplified model for the star formation history. Here we investigate the influence of the assumed star formation history of the Universe and the metallicity distribution of that star formation as well as some different binary evolution models on the WD AGWB. The paper is organised as follows. In Sect. 2 we present the model we use for the calculation of the AGWB as well as the different assumptions about binary evolution and star formation histories. In Sect. 3 we present the results of our calculations and we conclude with a discussion of the limitations of calculation, possible future work and the main conclusions (Sect. 4).
@@ 2.1 Calculating the WD AGWB
In order to calculate the GW flux, the population of WD binaries emitting GWs in a particular band in a redshift bin need to be simulated. To do so, two crucial ingredients are needed: a model for how binaries evolve from their formation as main sequence stars to the stage where they emit GWs plus a model for how many binaries were formed in the history of the Universe, the star formation rate density (SFRD, i.e. the amount of star formation per unit time per unit volume) over cosmic time. In SN24, the SFRD was approximated by a single function as given by Madau & Dickinson (2014). For the binary model, a single model at Solar metallicity, based on the SeBa population synthesis code (Portegies Zwart & Verbunt, 1996; Nelemans et al., 2001b; Toonen et al., 2012) was used. In the next sections we will introduce the variations we use in this study.
@@ 2.2 The Star Formation Rate Density
Figure 2: Total star formation rate density (SFRD) in $M_{\odot}$ Mpc-3 yr-1 versus redshift (z), for the six different SFRD models.
@@
With these models the metallicity evolution of the star formation can be followed. As an example we show in Fig. 1 the SRFD of the MZ19 model, split in six different metallicity bins, with bin centers ranging from Z = 0.0001 to Z = 0.03. The exact ranges are [0, 0.0005], (0.0005,0.003], (0.003,0.008], (0.008,0.015], (0.015,0.03], ($>$0.03). As expected, at high redshift the star formation is dominated by low metallicities, while in the lower redshift range the higher metallicities dominate. The SFRD split into the six metallicity bins for the LZ21 and HZ21 model are shown in Appendix A (the LZ19 and HZ19 plots are very similar but lower total due to the lack of starburst galaxies).
@@ 2.3 Binary population models
For each of the binary models we run the code for six different metallicities, to cover the bins in which we split up the SRFD above: Z = 0.0001, 0.001, 0.005, 0.01, 0.02, 0.03. For each run we simulate 250,000 binaries with primary masses between 0.8 and 11 $M_{\odot}$, distributed according to the Kroupa (2001) IMF, with a flat mass ratio distribution and initial separations flat in $\log a$. Taking into account the whole mass range (i.e. also those not simulated in detail) and a mass-dependent binary fraction (based on Moe & Di Stefano, 2017; van Haaften et al., 2013), this corresponds to a total mass of star formation of $3.4\times 10^{6}$ $M_{\odot}$.
For the formation of double WDs, one of the most important uncertainties is the effect of the two mass-transfer phases on the orbit. Because the stars typically start mass transfer when the donor has a deep convective envelope, the mass transfer is assumed to be unstable (e.g. Han 1998 but see Temmink et al., 2023). Earlier studies have suggested that the standard energy-balance ($\alpha$) common-envelope cannot be a good description of the first phase of mass transfer, but that a criterion based on angular momentum balance ($\gamma$) fits better (see Toonen et al., 2012; Nelemans & Tout, 2005; Nelemans et al., 2001b). In the classical common-envelope description, $\alpha$ nominally represents the efficiency with which orbital energy is used to unbind the envelope, i.e. with a strict upper bound on $\alpha$ of 1. However, the common-envelope phase is complicated and additional energy sources may be available, potentially leading to higher values. Indeed reconstruction of double WD evolution, suggests values as high as 4 (assuming a structure parameter $\lambda=0.5$, see e.g. Nelemans & Tout, 2005).
@@
In Fig. 3 we show the initial distribution of double WDs for this standard model as a function of initial GW frequency and chirp mass for the six different metallicities. Apart from the lowest metallicity, the bulk of the population is concentrated at low chirp mass and frequencies just above the line that indicates significant orbital evolution in a Hubble time. The distribution changes only slowly with metallicity. For the other models the distributions are shown in Appendix A. They all have similar average chirp masses around 0.45 $M_{\odot}$ with slightly higher values for the lowest metallicities.
@@ 3.1 The effect of metallicity
We first look at the influence of metallicity on the WD component of the AGWB in an artificial way, by simulating the AGWB assuming all star formation is at a particular metallicity. The result is shown in Fig. 4, where the green and the blue line represent the estimate for the BBH and BNS component and the dark red to yellow signals represent the WD component of the AGWB for the six different metallicities, respectively. The SFRD introduced in Madau & Dickinson (2014) is used for each of the metallicities. It is clear to see that a metallicity of 0.0001 leads to the lowest signal, whereas a metallicity of 0.005 leads to the highest signal. The other four metallicities are indistinguishable. The difference between the various signals, coming from the use of different metallicities, is approximately a factor of 1.9 at a frequency of 1 mHz. However, all six different metallicities lead to a WD component of the AGWB that is clearly dominant over the BBH and BNS components of the AGWB. Fig. 4 shows the results for a population model of $\gamma\alpha$ with $\alpha$ = 4. The three other population models are shown in Appendix A.
@@ 3.4 Estimation of the AGWB
Figure 8: As in Fig. 4. The red line represents the WD component of the AGWB for a choice of $\alpha\alpha$, $\alpha$ = 4, with Z = 0.02 and the SFRD of Madau & Dickinson (2014), which is the result from SN24. The light green line represents the WD component of the AGWB for a choice of $\gamma\alpha$, $\alpha$ = 4 with the MZ19 SFRD model. The light green band represents the uncertainty estimate.
@@
We show the final results of this work in Fig. 8, where the standard components are as in Fig. 4. The red line shows the results from the previous work as shown in SN24. In that work, a population model of $\alpha\alpha$ with $\alpha$ = 4 and a metallicity of 0.02, as well as the SFRD by Madau & Dickinson (2014) were used. The light green corresponds to the WD component of the AGWB for our standard model $\gamma\alpha$ with $\alpha$ = 4 and the MZ19 SFRD model. The light green band around the AGWB signal represents the uncertainties that arose in this work. We find these uncertainties by determining the upper and lower limits of the AGWB signal at all frequencies that come from the use of the different metallicity dependent SFRD models combined with all four binary population models. We find a strength of the WD component of the AGWB at 1 mHz of $\Omega_{\mathrm{WD}}(1\mathrm{mHz})=4.01_{-2.03}^{+5.63}\times 10^{-12}$. Just before the turnover, at a frequency of about 7 mHz, we find the following results: $\Omega_{\mathrm{WD}}(7\mathrm{mHz})=1.43_{-0.68}^{+1.95}\times 10^{-11}$. Looking at the factor between the upper and lower limits of the AGWB signal at 1 mHz and 7 mHz, respectively, we find a factor of approximately 5 difference between the upper and the lower limits.
@@ 4 Discussion and Conclusions
In conclusion, we investigated the influence of metallicity, different star formation estimates across cosmic time and different descriptions of the common-envelope phase on the level and shape of the WD AGWB. We conclude that the metallicity does not significantly impact the results, since only the very lowest metallicity (Z = 0.0001) produces a lower signal. However, the uncertainties in the (total amount) of star formation give rise to an uncertainty of about a factor of 5 in the level of the signal, while the different binary populations have a smaller effect on the level, but also influence the shape of the turnover of the signal. The uncertainty in the model suggests detection of the signal by a future GW detector would provide information about the star formation history and the binary properties at significant redshifts. The uncertainties are not large enough to challenge the conclusion that in the mHz regime the AGWB will be dominated by the WD population, significantly above the contribution of BHs.
@@ Acknowledgements.
We thank Seppe Staelens for his earlier work, discussions and help. G.N. is supported by the Dutch science foundation NWO.
@@ Appendix A Additional Figures
Figure 10: Total star formation rate density (SFRD) in $M_{\odot}$ Mpc-3 yr-1 versus redshift (z), for the six metallicity bins for the LZ21 SFRD model.
Figure 11: Total star formation rate density (SFRD) in $M_{\odot}$ Mpc-3 yr-1 versus redshift (z), for the six metallicity bins for the HZ21 SFRD model.

### v2

@@ Abstract
Context. The astrophysical gravitational wave background (AGWB) is a stochastic gravitational wave (GW) signal that is emitted by different populations of inspiralling binary systems containing compact objects throughout the Universe. In the frequency range between 10-4 and 10-1 Hz it will be detected by future space-based gravitational wave detectors like the Laser Interferometer Space Antenna (LISA). Recently, we concluded that the white dwarf (WD) contribution to the AGWB dominates over that of black holes (BHs) and neutron stars (NSs).
@@
Methods. We use the code developed before to determine the WD component of the AGWB. We use a metallicity dependent SFRD based on earlier work to construct five different SFRD models. We use four different population models that use different common-envelope treatment and six different metallicities for each model.
@@ 1 Introduction
In the visible Universe, there are a very large number of gravitational wave (GW) sources that consist of two stellar mass compact objects, white dwarfs (WDs), neutron stars (NSs) or black holes (BHs). With a GW detector typically only a (tiny) selection of these sources can be individually detected, and many cannot. The Laser Interferometer Space Antenna (LISA) mission (Amaro-Seoane et al., 2017), for instance can individually detect a large number of binaries (mostly double WD) in the Milky Way (e.g. Lipunov et al., 1987; Evans et al., 1987; Hils et al., 1990; Nelemans et al., 2001a; Ruiter et al., 2010; Yu & Jeffery, 2013; Lamberts et al., 2019; Korol et al., 2022; Temmink et al., 2023). The remaining (millions) of binaries in the Milky Way form a collective stochastic foreground of Galactic sources that is highly anisotropic (e.g. Nissanke et al., 2012; Karnesis et al., 2021). However, there is a huge number of unresolved compact binary sources in the extragalactic Universe that form a stochastic isotropic signal (e.g. Farmer & Phinney, 2003). This Astrophysical GW Background (AGWB) , the background signal in GW solely made up of astrophysical sources, encodes the combined information about the different source populations (see, e.g. Schneider et al., 2001; Farmer & Phinney, 2003; Kowalska-Leszczynska et al., 2015; Amaro-Seoane et al., 2023). It is a broadband signal (Phinney, 2001) that extends from the high-frequency band covered by the LIGO/Virgo/KAGRA detectors (LIGO Scientific Collaboration et al., 2015; Acernese et al., 2015; Aso et al., 2013) to the mHz band covered by future space GW detectors such as LISA (Amaro-Seoane et al., 2017), Taiji (Luo et al., 2021) or TianQin (Luo et al., 2016). Its amplitude gives an estimate of the total number of sources, weighted with their redshift distribution, while its shape potentially encodes details of the population, something we study in more detail in this paper. Since it is not likely that in particular the WD populations can be probed in any other way, this provides a unique probe of binary evolution at high redshift.
@@
The AGWB is not only interesting in itself as probe of the large population of binaries across cosmic time that cannot be detected individually, but also competes with potential other stochastic GW backgrounds that have an origin in the Early Universe and probe fundamental physics (see e.g. Caprini et al., 2009; Binétruy et al., 2012; Auclair et al., 2023). While LISA will be able to detect the AGWB, one of the main objectives is to be able to detect the cosmological GWB (caused by e.g. cosmic strings (Caprini & Figueroa, 2018; Auclair et al., 2023)). Understanding of the AGWB will therefore be of considerable importance to achieve the LISA objective to investigate the cosmological GWB (Lehoucq et al., 2023).
@@
In the LISA band, the AGWB due to BH mergers could be detected (e.g. Dvorkin et al., 2016; Bavera et al., 2022; Babak et al., 2023; Lehoucq et al., 2023), but Staelens & Nelemans (2024) recently showed that the AGWB due to WD binaries (studied earlier by Farmer & Phinney, 2003) likely dominates over that due to BHs. The AGWB due to NSs is expected to be significantly below even that of BHs because of the relative small number of NS mergers compared to WD mergers and their relative low mass compared to BHs (e.g. Babak et al., 2023; Abbott et al., 2023). However, Staelens & Nelemans (2024) only considered a single model for the double WD population in the Universe and a simplified model for the star formation history. Yet, we know that the metallicity of the progenitor stars may have a significant impact on the compact binary formation (e.g. Giacobbo & Mapelli, 2018; Chruslinska et al., 2018; Neijssel et al., 2019) and the metallicity of star formation is different at different times in the Universe. In addition, the star formation history of the Universe is uncertain (e.g. Chruślińska & Nelemans, 2019). Here, we investigate the influence of the assumed star formation history of the Universe and the metallicity distribution of that star formation as well as some different binary evolution models on the WD AGWB.
The paper is organised as follows. In Sect. 2 we present the model we use for the calculation of the AGWB as well as the different assumptions about binary evolution and star formation histories. In Sect. 3 we present the results of our calculations and we conclude with a discussion of the limitations of calculation, possible future work and the main conclusions (Sect. 4).
@@ 2.1 Calculating the WD AGWB
In order to calculate the GW flux, the population of WD binaries emitting GWs in a particular band in a redshift bin needs to be simulated. To do so, two crucial ingredients are needed: a model for how binaries evolve from their formation as main sequence stars to the stage where they emit GWs plus a model for how many binaries were formed in the history of the Universe, the star formation rate density (SFRD, i.e. the amount of star formation per unit time per unit volume) over cosmic time. In SN24, the SFRD was approximated by a single function as given by Madau & Dickinson (2014). For the binary model, a single model at Solar metallicity, based on the SeBa population synthesis code (Portegies Zwart & Verbunt, 1996; Nelemans et al., 2001b; Toonen et al., 2012) was used. In the next sections we will introduce the variations we use in this study.
@@ 2.2 The star formation rate density
Figure 2: Total star formation rate density (SFRD) in $M_{\odot}$ Mpc-3 yr-1 versus redshift (z), for the six different SFRD models. The red line represents the Madau & Dickinson (2014) model used in SN24. The other five lines represent five of the SFRD models introduced in Chruślińska & Nelemans (2019); Chruślińska et al. (2020, 2021). The bottom three models are the low- intermediate- and high-metallicity extremes without starbursts and the top two the extreme models including starbursts (see text for more details).
@@
With these models the metallicity evolution of the star formation can be followed. As an example we show in Fig. 1 the SRFD of the MZ19 model, split in six different metallicity bins, with bin centers ranging from Z = 0.0001 to Z = 0.03. The exact ranges are [0, 0.0005], (0.0005, 0.003], (0.003, 0.008], (0.008, 0.015], (0.015, 0.03], ($>$0.03). As expected, at high redshift the star formation is dominated by low metallicities, while in the lower redshift range the higher metallicities dominate. The SFRD split into the six metallicity bins for the LZ21 and HZ21 model are shown in Appendix A (the LZ19 and HZ19 plots are very similar but a lower total due to the lack of starburst galaxies).
@@ 2.3 Binary population models
For each of the binary models we run the code for six different metallicities, to cover the bins in which we split up the SRFD above: Z = 0.0001, 0.001, 0.005, 0.01, 0.02, 0.03. For each run we simulate 250,000 binaries with primary masses between 0.8 and 11 $M_{\odot}$, distributed according to the Kroupa (2001) initial mass function, with a flat mass ratio distribution and initial separations flat in $\log a$. Taking into account the whole mass range (i.e. also those not simulated in detail) and a mass-dependent binary fraction (based on Moe & Di Stefano, 2017; van Haaften et al., 2013), this corresponds to a total mass of star formation of $3.4\times 10^{6}$ $M_{\odot}$.
For the formation of double WDs, one of the most important uncertainties is the effect of the two mass-transfer phases on the orbit. Because the stars typically start mass transfer when the donor has a deep convective envelope, the mass transfer is assumed to be unstable (e.g. Han 1998 but see Temmink et al., 2023). Earlier studies have suggested that the standard energy-balance ($\alpha$) common-envelope cannot be a good description of the first phase of mass transfer since it inevitably leads to double WDs in which the last formed WD has a lower mass, which is not observed, but that a criterion based on angular momentum balance ($\gamma$) fits better (see Toonen et al., 2012; Nelemans & Tout, 2005; Nelemans et al., 2001b). Some works have suggested the first mass transfer is in fact stable (Woods et al., 2012; Li et al., 2023), but that assumption leads to the opposite situation in which the last formed WD is the more massive one, which is also not always the case. In the classical common-envelope description, $\alpha$ nominally represents the efficiency with which orbital energy is used to unbind the envelope, i.e. with a strict upper bound on $\alpha$ of 1. However, the common-envelope phase is complicated and additional energy sources may be available, potentially leading to higher values. Indeed reconstruction of double WD evolution, suggests values as high as 4 (assuming a structure parameter $\lambda=0.5$, see e.g. Nelemans & Tout, 2005).
@@
In Fig. 3 we show the initial distribution of double WDs for this standard model as a function of initial GW frequency and chirp mass for the six different metallicities. Apart from the lowest metallicity, the bulk of the population is concentrated at low chirp mass and frequencies just above the line that indicates significant orbital evolution in a Hubble time. The lowest metallicity looks different, because at low metallicity the stars are more compact leading to more massive cores when stars interact. This leads to higher chirp masses, but more importantly, after the first mass transfer, to tighter orbits that in the second mass transfer lead to merger of the system or, in case the more compact stars makes that the first mass transfer avoids the first giant branch, to wider systems that lead to wider final orbits. The distribution changes only slowly with metallicity. For the other models the distributions are shown in Appendix A. They all have similar average chirp masses around 0.45 $M_{\odot}$ with slightly higher values for the lowest metallicities.
@@ 3.1 The effect of metallicity
We first look at the influence of metallicity on the WD component of the AGWB in an artificial way, by simulating the AGWB assuming all star formation is at a particular metallicity. The result is shown in Fig. 4, where the green and the blue line represent the estimate for the binary black hole (BBH) and binary neutron star (BNS) component (see Abbott et al. 2023) and the dark red to yellow signals represent the WD component of the AGWB for the six different metallicities, respectively. The SFRD introduced in Madau & Dickinson (2014) is used for each of the metallicities. It is clear to see that a metallicity of 0.0001 leads to the lowest signal, whereas a metallicity of 0.005 leads to the highest signal. The other four metallicities are indistinguishable. The difference between the various signals, coming from the use of different metallicities, is approximately a factor of 1.9 at a frequency of 1 mHz. However, all six different metallicities lead to a WD component of the AGWB that is clearly dominant over the BBH and BNS components of the AGWB. Fig. 4 shows the results for a population model of $\gamma\alpha$ with $\alpha$ = 4. The three other population models are shown in Appendix A.
@@ 3.4 Estimation of the AGWB
Figure 8: As in Fig. 4. The red line represents the WD component of the AGWB for a choice of $\alpha\alpha$, $\alpha$ = 4, with Z = 0.02 and the SFRD of Madau & Dickinson (2014), which is the result from SN24. The light green line represents the WD component of the AGWB for a choice of $\gamma\alpha$, $\alpha$ = 4 with the MZ19 SFRD model. The light green band represents the uncertainty estimate. The solid grey line shows the normal LISA sensitivity (i.e. not integrated over time and frequency).
@@
We show the final results of this work in Fig. 8, where the standard components are as in Fig. 4. The red line shows the results from the previous work as shown in SN24. In that work, a population model of $\alpha\alpha$ with $\alpha$ = 4 and a metallicity of 0.02, as well as the SFRD by Madau & Dickinson (2014) were used. The light green corresponds to the WD component of the AGWB for our standard model $\gamma\alpha$ with $\alpha$ = 4 and the MZ19 SFRD model. The light green band around the AGWB signal represents the uncertainties that arose in this work. We find these uncertainties by determining the upper and lower limits of the AGWB signal at all frequencies that come from the use of the different metallicity dependent SFRD models combined with all four binary population models. We find a strength of the WD component of the AGWB at 1 mHz of $\Omega_{\mathrm{WD}}(1\mathrm{mHz})=4.01_{-2.03}^{+5.63}\times 10^{-12}$. Just before the turnover, at a frequency of about 7 mHz, we find the following results: $\Omega_{\mathrm{WD}}(7\mathrm{mHz})=1.43_{-0.68}^{+1.95}\times 10^{-11}$. Looking at the factor between the upper and lower limits of the AGWB signal at 1 mHz and 7 mHz, respectively, we find a factor of approximately 5 difference between the upper and the lower limits. We also show the normal LISA sensitivity, which is above the WD AGWB, to make clear that the WD AGWB does not significantly impact the other LISA science.
@@ 4 Discussion and conclusions
In conclusion, we investigated the influence of metallicity, different star formation estimates across cosmic time and different descriptions of the common-envelope phase on the level and shape of the WD AGWB. We conclude that the metallicity does not significantly impact the results, since only the very lowest metallicity (Z = 0.0001) produces a lower signal. However, the uncertainties in the (total amount) of star formation give rise to an significant uncertainty in the level of the signal, while the different binary populations have a smaller effect on the level, but also influence the shape of the turnover of the signal. The total uncertainty on the level of the signal is about a factor of 5. The uncertainty in the model suggests detection of the signal by a future GW detector would provide information about the star formation history and the binary properties at significant redshifts. The uncertainties are not large enough to challenge the conclusion that in the mHz regime the AGWB will be dominated by the WD population, significantly above the contribution of BHs.
@@ Acknowledgements.
We thank Seppe Staelens for his earlier work, discussions and help and the referee for valuable suggestions and comments. G.N. is supported by the Dutch science foundation NWO.
@@ Appendix A Additional Figures
In this appendix we provide additional figures that give more details of the different models.
Firstly, in the main paper we present the split of the SFRD over six different metallicity bins for the MZ19 model in Fig. 1. Figs. 10 and Fig. 11 give the same for the LZ21 and the HZ21 SFRD models. The LZ19 and HZ19 SFRD model are left out, since these are very similar except for the normalisation due to the inclusion of starburst galaxies in the new models.
Figure 10: Total SFRD versus redshift (z), for the six metallicity bins for the LZ21 SFRD model.
Figure 11: Total SFRD versus redshift (z), for the six metallicity bins for the HZ21 SFRD model.
In the main text we presented the results for our default binary evolution model ($\gamma\alpha$, with $\alpha$ = 4). Figs. 15 to 17 show the initial GW frequency and chirp mass for the different metallicities for the population models $\alpha\alpha$, with $\alpha$ = 1, $\alpha\alpha$, with $\alpha$ = 4 and $\gamma\alpha$, with $\alpha$ = 1, respectively. As for the standard model, most metallicities show very similar structure with gradual changes, mostly towards lower GW frequency and higher chirp mass for lower metallicity, with the lowest metallicity model often being quite different. As for the standard model, this is mostly due to the fact that stars are more compact at low metallicity.
The resulting AGWB spectra for the hypothetical cases that all star formation would happen at a single metallicity are shown in Fig. 12 to 14. These are the same as Fig. 4, except for the other population models. For the $\alpha\alpha$, with $\alpha$ = 1 model the shorter periods after the common envelope lead to an increasing number of sources with higher chirp mass towards lower metallicity, making the lower metallicity AGWBs higher than those of higher metallicity. This is contrary to in our standard model, but the overall AGWB is lower in these models. The $\alpha\alpha$, with $\alpha$ = 4 model shows similar behaviour to the standard model, with somewhat higher overall levels. Finally, in the $\gamma\alpha$, with $\alpha$ = 1 model, the increase in chrip mass towards lower metallicity is compensated with decrease in over all number, leading to very similar overall AGWB levels for all metallicities.